In the process of being felled, a tree is pushed horizontally to avoid hitting other trees as it falls. It feels a net force with a magnitude of 20,000 N as it falls. If the horizontal force on the tree is 2900 N, what is the weight of the tree?


1
Expert's answer
2021-01-19T11:59:10-0500


The net force is a vector sum of the two other forces: vertical (weight) and horizontal (pull):


"F_\text{net}^2=W^2+P^2,\\\space\\
W=\sqrt{F_\text{net}^2-P^2}=19789\text{ N}."